A jet airliner moving initially at 3.50 ✕ 102 mi/h due east enters a region where the wind is blowing at 1.00 ✕ 102 mi/h in a direction 35.0° north of east. (Let the x-direction be eastward and the y-direction be northward.)

A. Find the components of the velocity of the jet airliner relative to the air, vJA.

vJA,x = _____mi/h

vJA,y = _____mi/h

B. Find the components of the velocity of the air relative to Earth, vAE.

vAE,x = ____mi/h

vAE,y = ____mi/h

C. Write an equation analogous to vAB = vAE − vBE for the velocities vJA, vAE, and vJE . (In the given equation, vBE is the velocity of observer B relative to observer E. Object A travels with velocity vAB relative to observer B, and velocity vAE relative to observer E.)

(option a) vJA = vJE + vAE

(option B) vJA = vAE − vJE

(option C) vJA = vAE + 2vJE

(option d) vJA = vJE − vAE

D. What are the speed and direction of the aircraft relative to the ground?

magnitude _____mi/h

direction _______° north of east
Answer #1

(A) The components of the velocity of jet airliner relative to an air which is given as :

On x-axis, we have

vJA,x = 350 mph

On y-axis, we have

vJA,y = 0 mph

(b) The components of the velocity of an air relative to Earth which is given as :

On x-axis, we have

vAE,x = vJA,x + vJE,x

vAE,x = (350 mph) + (100 mph) cos 350

vAE,x = [(350 mph) + (81.9 mph)]

vAE,x = 431.9 mph

On y-axis, we have

vAE,y = vJA,y + vJE,y

vAE,y = (0 mph) + (100 mph) sin 350

vAE,y = [(0 mph) + (57.3 mph)]

vAE,y = 57.3 mph

(c) Write an equation analogous to vAB = vAE − vBE for the velocities vJA, vAE, and vJE .

where, vBE = velocity of observer B relative to observer E

vAB = object A travels with velocity relative to observer B

vAE = object A travels with velocity relative to observer E

(option B) : vJA = vAE − vJE

(d) The speed and direction of an aircraft relative to the ground which will be given as :

We know that, vAE = vAB + vBE

vAE = \sqrt{}vAE,x2 + vAE,y2

vAE = \sqrt{}(431.9 mph)2\hat{}+ (57.3 mph)2

vAE = 435.6 mph

Using a trigonometric identity, we have

\theta = tan-1 [(57.3 mph) / (431.9 mph)]

\theta = 7.55 degree

{ north of east }